Circular No. 7530

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000ew IN NGC 3810 T. Puckett and A. Langoussis, Mountain Town, GA, report the dis-

covery of an apparent supernova (mag 14.9) on an unfiltered CCD frame (limiting mag 18.5) taken with the Puckett Observatory 0.30-m automated supernova patrol telescope on Nov. 28.48 UT. The candidate is located at α = 11^{h}40^{m}58^{s}.52, δ = +11^{o}27′55′′.9 (equinox 2000.0), which is 1′′.9 west and 20′′.7 south of the center of NGC 3810. The new object was also present and confirmed (measured end figures 58^{s}.56, 56′′.4) on unfiltered CCD frames taken with a 0.6-m reflector by G. J. Garradd, Biosphere 2 Observatory, on Nov. 29.39. The candidate does not appear on unfiltered images taken by Puckett during Feb. 8–May 8, and it does not appear on Palomar Sky Survey images taken on 1989 Apr. 4 and 1991 Apr. 17 (each limiting mag ∼ 21.0). SN 1997dq was located in this same galaxy (cf. IAUC 6770).

SUPERNOVA 2000ex IN ESO 419-G3 A. B. Aazami and W. D. Li, University of California, Berkeley (UCB),

on behalf of LOTOSS (cf. IAUC 7514), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova on unfiltered images taken on Nov. 26.4 (mag ∼ 17.6) and 27.3 UT (mag ∼ 17.8). The new object is located at α = 3^{h}42^{m}11^{s}.71, δ = −27^{o}51′48′′.0 (equinox 2000.0), which is 7′′.6 east and 3′′.4 north of the nucleus of ESO 419-G3 = MCG –05-09-022. A KAIT image of the same field on Nov. 17.4 also showed the supernova at mag ∼ 17.8, while an image on Oct. 19.4 showed nothing at this position (limiting mag ∼ 19.0). SN 1978C was located in this same galaxy (cf. IAUC 3254).

T. Matheson, S. Jha, P. Challis and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 2000ex, obtained by K. Rines on Nov. 28.28 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type II supernova ∼ 1 week after maximum light. The spectrum consists of a blue continuum with superposed P-Cyg lines of hydrogen and Ca II. Narrow Hα emission superposed on the spectrum of the supernova indicates a recession velocity of 4090 km/s for the host galaxy. The expansion velocity derived from the minimum of the Hβ line is ∼ 5000 km/s.

A. V. Filippenko and R. Chornock, UCB, report that inspection of noisy CCD spectra (range 330–760 nm) obtained on Nov. 29 with the Shane 3-m reflector at Lick Observatory reveals that the object is a normal type II supernova.

2000 November 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden